OpenStudy (anonymous):

Use properties of logarithms to rewrite expression. log_t⁡∛((x^3 y^5)/(tz^2 ))

OpenStudy (anonymous):

log(a+b) = loga+logb log(a-b) = loga/logb logx^n = nlogx

OpenStudy (isaiah.feynman):

"log(a+b) = loga+logb" That is wrong. The correct rule is log(a*b)=loga+logb

OpenStudy (anonymous):

Isaiah.Feynman do you think you can help me figure this out? I understand the rules but this problem is a lot harder than others that I have solved.

OpenStudy (anonymous):

oops, sorry I mean log(a*b) = loga+logb and log(a/b) = loga-logb

OpenStudy (anonymous):

Okay I understand those rules, I just have trouble with the square roots and exponents. Do you think you can help me figure this problem out ECE?

OpenStudy (anonymous):

Can you draw out your problem? I can't read it in the form you wrote sorry

OpenStudy (jdoe0001):

\(\huge log_t\sqrt[3]{\cfrac{x^3y^5}{tz^2}}\quad ?\)

OpenStudy (anonymous):

yes it is